ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: El Kinani, E. H., Département de Physique, L. P. T , Faculté des Sciences Mohammed V, B. P 1014, Rabat , Morocco q-Derivation, Sine algebra and Ut(sl(2)) We review the algebra of q-pseudo-differential operators (q − ψ DO) and certain of its properties. Based on the fact that w(∞) and Ut(sl(2)) can be constructed from the trigonometric sine algebra(FFZ algebra), we propose the embedding of w(∞) and Ut(sl(2)) in (q − ψ DO). The correspondance with 2-D fractional supercharge is examined .